# Pentagram maps over rings, Grassmannians, and skewers

### v2

@@ 5.1 Difference operators with coefficients in a ring
The ring of (left) difference operators over $R$ is a subring $R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by operators of left multiplication by elements of $R^{\mathbb{Z}}$, along with the left shift operator $T\in\mathrm{End}_{R}(R^{\mathbb{Z}})$, given by $(Ta)_{i}:=a_{i+1}$. In other words, a difference operator is an endomorphism $\mathcal{D}:R^{\mathbb{Z}}\to R^{\mathbb{Z}}$ of right $R$-modules given by
@@
where $d\geq 0$ is an integer, and $\alpha_{1},\dots,\alpha_{d}\in R^{\mathbb{Z}}$. We will denote the entries of the bi-infinite sequence $\alpha_{i}$ by $\alpha_{ij}$. For any $a\in R^{\mathbb{Z}}$, the operator $\mathcal{D}$ acts on $a$ as
@@
Assuming $\alpha_{d}\neq 0$, we call $d$ the degree of the difference operator $\mathcal{D}$.
More generally, one can consider the subring $R^{\mathbb{Z}}[T,T^{-1}]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by operators of left multiplication by elements of $R^{\mathbb{Z}}$, along with $T$ and $T^{-1}$, i.e., allow for difference operators of the form
@@
One can also define right difference operators: they are endomorphisms of $R^{\mathbb{Z}}$ viewed as a left $R$-module. A right difference operator reads
$(\mathcal{D}a)_{j}=\sum_{i=0}^{d}a_{i+j}\alpha_{ij}.$
However, for the remainder of the paper, all difference operators are assumed to be left difference operators, unless stated otherwise.
@@
Let $\mathcal{D}=\sum_{i=0}^{d}\alpha_{i}T^{i}\in\mathbb{R}^{\mathbb{Z}}[T]$ be a difference operator. If $\alpha_{0}$ and $\alpha_{d}$ are sequences of units in $R$ (i.e., are units in $R^{\mathbb{Z}}$), we call $\mathcal{D}$ a properly bounded difference operator.
@@
If we have a difference operator $\mathcal{D}\in\mathbb{R}^{\mathbb{Z}}[T]$ which is not properly bounded, there is no bound on the “size” of the kernel in terms of the degree. In particular, the kernel does not have to be free and/or of finite rank. Nevertheless, we have the following.
@@
Let $\mathcal{D}\in\mathbb{R}^{\mathbb{Z}}[T]$, $\mathcal{D}\neq 0$ be a difference operator of degree less than $d$. Then, for any properly bounded difference operator $\mathcal{D}^{\prime}\in\mathbb{R}^{\mathbb{Z}}[T]$ of degree $d$, there exists $a\in\mathrm{Ker}\,\mathcal{D}^{\prime}$ such that $a\notin\mathrm{Ker}\,\mathcal{D}$.
@@
Since $\mathcal{D}\neq 0$, there is $i\in\mathbb{Z}$ such that the condition $\mathcal{D}a=0$, where $a\in R^{\mathbb{Z}}$, imposes a non-trivial relation on $d$ consecutive entries $a_{i},\dots,a_{i+d-1}$ of $a$. On other hand, the map $\mathrm{Ker}\,\mathcal{D}^{\prime}\to R^{d}$ given by $a\mapsto(a_{i},\dots,a_{i+d-1})$ is a bijection, so we can find an element of $\mathrm{Ker}\,\mathcal{D}^{\prime}$ whose entries do not satisfy the relation forced by $\mathcal{D}a=0$. ∎
Now, let $M$ be a free finite-rank $R$-module. Then $R^{\mathbb{Z}}\otimes_{R}M\simeq M^{\mathbb{Z}}$, as left $R^{\mathbb{Z}}$-modules (where $M^{\mathbb{Z}}$ consists of bi-infinite sequences valued in $M$). Any difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ acts on $M^{\mathbb{Z}}=R^{\mathbb{Z}}\otimes_{R}M$ by $\mathcal{D}(a\otimes v)=(\mathcal{D}a)\otimes v$, where $a\in R^{\mathbb{Z}}$, $v\in M$. Put differently, for $\mathcal{D}=\sum_{i=0}^{d}\alpha_{i}T^{i}$ and $v\in M^{\mathbb{Z}}$, we have
@@
($\ v_{i,1}&\dots&v_{i,m}\\ \vdots&&\vdots\end{array}\right)$ (5.1)
@@
Suppose $\mathcal{D}\in\mathbb{R}^{\mathbb{Z}}[T]$ is a difference operator, and $v\in M^{\mathbb{Z}}$, where $M$ is a free finite-rank left $R$-module. Then $\mathcal{D}v=0$ if and only if $\mathrm{Span}\,v\subset\mathrm{Ker}\,\,\mathcal{D}$.
@@
which is 00 if and only if $\mathcal{D}(a_{i}){\,=0}$ for all $i$, which is the same as to say $\mathrm{Span}\,v\subset\mathrm{Ker}\,\,\mathcal{D}$. ∎
The following proposition is a difference counterpart of the following standard result from the theory of ordinary differential equations: for functions $f_{1}(x),\dots,f_{n}(x)$ belonging to a kernel of the differential operator
@@
Let $R$ be a stably finite ring. Suppose $\mathcal{D}\in\mathbb{R}^{\mathbb{Z}}[T]$ is a properly bounded difference operator of degree $d$, and $v\in M^{\mathbb{Z}}$, where $M$ is a free left $R$-module of rank $d$. Suppose that $\mathcal{D}v=0$. Then the following conditions are equivalent:
@@
$V_{k}:=\left(\begin{array}[]{ccc}v_{k,1}&\dots&v_{k,d}\\ \vdots&&\vdots\\ v_{k+d-1,1}&\dots&v_{k+d-1,d}\end{array}\right)\in\text{M}_{d}(R)$
where $v_{i,j}\in R$ are coordinates of $v_{i}$ relative to the basis $e_{1},\dots,e_{d}$. The rows of the matrix $V_{k}$ are coordinate expressions for $v_{k},\dots,v_{k+d-1}\in M$, so $V_{k}$ is left invertible if and only if $v_{k},\dots,v_{k+d-1}$ span $M$. At the same time, the matrix $V_{k}$ consists of $d$ consecutive rows of the infinite matrix (5.1) whose columns span the module $\mathrm{Span}\,v$. Since taking $d$ consecutive entries gives an isomorphism $\mathrm{Ker}\,\,\mathcal{D}\simeq R^{d}$, the columns of the matrix $V_{k}$ span $R^{d}$ if and only if the columns of the matrix (5.1) span $\mathrm{Ker}\,\,\mathcal{D}$, which is equivalent to saying that $\mathrm{Span}\,v=\mathrm{Ker}\,\,\mathcal{D}$. To sum up, $v_{k},\dots,v_{k+d-1}$ span $M$ is and only if the matrix $V_{k}$ is left-invertible, while $\mathrm{Span}\,v=\mathrm{Ker}\,\,\mathcal{D}$ if and only if $V_{k}$ is right-invertible. Since $R$ is stably finite, those conditions are equivalent. ∎
@@ 5.2 The correspondence between difference operators and polygons
Note that the ring $\mathbb{R}^{\mathbb{Z}}[T]$ of difference operators over $R$ contains a subring $R^{\mathbb{Z}}$ of bi-infinite sequences valued in $R$: such sequences are just degree zero difference operators. In particular, the group of units in $R^{\mathbb{Z}}$ (which consists of bi-infinite sequences of units in $R$) acts on $\mathbb{R}^{\mathbb{Z}}[T]$ in two different ways: by left multiplication, and by right multiplication.
@@
By the set of projective equivalence classes of polygons in the left projective space $\mathbb{P}^{d-1}(R)$, we mean the set polygons in projectivizations of all rank $d$ free left $R$-modules, considered up to projective equivalence.
Similarly, there is a natural one-to-one correspondence between the set of projective equivalence classes of polygons in the right projective space $\mathbb{P}^{d-1}(R)$, and the set of properly bounded right difference operators of degree $d$ with coefficients in $R$, up to multiplication on the left and right by bi-infinite sequences of units in $R$. This can be proved analogously to the left case, or deduced from the left case by viewing right $R$-modules as left modules over the opposite ring.
In the case when the base ring $R$ is a field, the correspondence between polygons and difference operators is well-known. In particular, [8] gives a proof of Theorem 5.7 in the case $R=\mathbb{R}$.
@@
Let us denote the space of all degree $d$ properly bounded difference operators over $R$ as $R^{\mathbb{Z}}[T]_{d}^{pb}$ and the space of polygons in $\mathbb{P}^{d-1}(R)$ up to projective equivalence as $\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Our first aim to build a map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
Let $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$. Then, by Proposition 5.3, we know that $\mathrm{Ker}\,\,\mathcal{D}\subset R^{\mathbb{Z}}$ is a rank $d$ free right $R$-module. Thus, $(\mathrm{Ker}\,{\mathcal{D}})^{*}=\mathrm{Hom}_{R}(\mathrm{Ker}\,{\mathcal{D}},R)$ is a free rank $d$ left $R$-module. For any $i\in\mathbb{Z}$, let $\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,{\mathcal{D}})^{*}$ be defined by $\varepsilon_{\mathcal{D},i}(a):=a_{i}$ for any $a\in\mathrm{Ker}\,{\mathcal{D}}$. Then $\varepsilon_{\mathcal{D}}\in((\mathrm{Ker}\,{\mathcal{D}})^{*})^{\mathbb{Z}}=R^{\mathbb{Z}}\otimes_{R}(\mathrm{Ker}\,{\mathcal{D}})^{*}$.
@@
We have $\mathrm{Span}\,\varepsilon_{\mathcal{D}}=\mathrm{Ker}\,\mathcal{D}$. In particular, $\mathcal{D}\varepsilon_{\mathcal{D}}=0$.
@@
For any $i\in\mathbb{Z}$, the list $\varepsilon_{\mathcal{D},i},\dots,\varepsilon_{\mathcal{D},i+d-1}$ is a basis of $(\mathrm{Ker}\,\,\mathcal{D})^{*}$.
@@
It follows that the sequence of points $P_{i}=R\varepsilon_{\mathcal{D},i}$ is a polygon in $\mathbb{P}((\mathrm{Ker}\,\,\mathcal{D})^{*})$. Thus, every properly bounded difference operator gives rise to a polygon in the projectivization of the dual to its kernel. This gives a map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
Denote the collection of properly bounded, degree $d$ left difference operators up to multiplication on the left and right by bi-infinite sequences of units as ${R^{\mathbb{Z}}[T]_{d}^{pb}}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$.
@@
The map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ factors through a map $\varphi:{R^{\mathbb{Z}}[T]_{d}^{pb}}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
@@
We need to check that if we take a properly bounded left difference operator $\mathcal{D}$, and multiply it on the left or right by a bi-infinite sequence of units in $R$, then the image of the resulting operator under $\bar{\varphi}$ is in the same projective equivalence class as the image of $\mathcal{D}$. Consider $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ and a bi-infinite sequence $c\in(\mathbb{R}^{\mathbb{Z}})^{*}$ of units in $R$. Note that multiplication on the left $\mathcal{D}\mapsto c\mathcal{D}$ does not change the kernel and hence the image under $\bar{\varphi}$. So, it suffices to establish projective equivalence of the polygons $\bar{\varphi}(\mathcal{D}c)$ and $\bar{\varphi}(\mathcal{D})$. Clearly, left multiplication by $c$ gives a right $R$-module isomorphism $\lambda_{c}\colon\mathrm{Ker}\,\mathcal{D}c\to\mathrm{Ker}\,\mathcal{D}$. Consider the associated left module isomorphism $\lambda_{c}^{*}\colon(\mathrm{Ker}\,\mathcal{D}c)^{*}\to(\mathrm{Ker}\,\mathcal{D})^{*}$. As before, let $\varepsilon_{\mathcal{D}c,i}\in(\mathrm{Ker}\,\mathcal{D}c)^{*}$ be defined by $\varepsilon_{\mathcal{D}c,i}(a)=a_{i}$, and let $a\in\mathrm{Ker}\,\mathcal{D}$. Then
@@
Suppose that $\mathcal{D},\mathcal{D}^{\prime}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ are properly bounded (left) difference operators of degree $d$. Assume also that $\mathrm{Ker}\,\mathcal{D}=\mathrm{Ker}\,\mathcal{D}^{\prime}$. Then $\mathcal{D}^{\prime}=c^{\prime}\mathcal{D}$ for some $c^{\prime}\in(R^{\mathbb{Z}})^{*}$.
@@
Since $\mathcal{D}^{\prime}$ and $\mathcal{D}$ are properly bounded of the same degree $d$, there is a sequence of units $c^{\prime}\in(R^{\mathbb{Z}})^{*}$ such that $\mathcal{R}:=\mathcal{D}^{\prime}-c^{\prime}\mathcal{D}$ has degree at most $d-1$. Note that since $\mathrm{Ker}\,\mathcal{D}^{\prime}=\mathrm{Ker}\,\mathcal{D}$, we have $\mathrm{Ker}\,\mathcal{R}\supset\mathrm{Ker}\,\mathcal{D}$. So, by Proposition 5.4, it follows that $\mathcal{R}=0$, meaning that $\mathcal{D}^{\prime}=c^{\prime}\mathcal{D}$, as desired. ∎
@@
The map $\varphi:{R^{\mathbb{Z}}[T]_{d}^{pb}}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ is injective.
@@
Consider two properly bounded left difference operators $\mathcal{D},\mathcal{D}^{\prime}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ such that $\overline{\varphi}(\mathcal{D})=\overline{\varphi}(\mathcal{D}^{\prime})$. We need to show that $\mathcal{D}$ and $\mathcal{D}^{\prime}$ belong to the same class in the quotient ${R^{\mathbb{Z}}[T]_{d}^{pb}}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$.
@@
Now, to complete the proof of Theorem 5.7, it suffices to show that the mapping $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ is surjective. In other words, given a polygon $(P_{i})$ in $\mathbb{P}(M)$. where $M$ is a rank $d$ free left $R$-module, we need to find a difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ such that the associated polygon in $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})$ is projectively equivalent to $(P_{i})$.

### v3

@@ 5.1 Difference operators with coefficients in a ring
The ring of (left) difference operators over $R$ is a subring $R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by operators of left multiplication by elements of $R^{\mathbb{Z}}$, along with the shift operator $T\in\mathrm{End}_{R}(R^{\mathbb{Z}})$, given by $(T{a})_{i}:=a_{i+1}$. In other words, a difference operator is an endomorphism $\mathcal{D}:R^{\mathbb{Z}}\to R^{\mathbb{Z}}$ of right $R$-modules given by
@@
where $d\geq 0$ is an integer, and $\alpha_{1},\dots,\alpha_{d}\in R^{\mathbb{Z}}$. Assuming $\alpha_{d}\neq 0$, we call $d$ the degree of the difference operator $\mathcal{D}$. We will denote the entries of the bi-infinite sequence $\alpha_{i}$ by $\alpha_{ij}$. For any $a\in R^{\mathbb{Z}}$, the operator $\mathcal{D}$ acts on $a$ as
@@
More generally, one can consider the subring $R^{\mathbb{Z}}[T,T^{-1}]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by operators of left multiplication by elements of $R^{\mathbb{Z}}$, along with $T$ and $T^{-1}$, that is, allow for difference operators of the form
@@
Let $\mathcal{D}=\sum_{i=0}^{d}\alpha_{i}T^{i}\in R^{\mathbb{Z}}[T]$ be a difference operator. If $\alpha_{0}$ and $\alpha_{d}$ are sequences of left invertible elements in $R$ (i.e., are left invertible in $R^{\mathbb{Z}}$), we call $\mathcal{D}$ a properly bounded difference operator.
In the next section, we will establish a correspondence between properly bounded difference operators and polygons. The remaining part of the current section contains various technical results on difference operators which will be used to establish that correspondence.
To a properly difference operator $\mathcal{D}$ of degree $d$, we will associate a polygon in the projectivization of $(\mathrm{Ker}\,\mathcal{D})^{*}$. The latter is a (d1)(d-1)-dimensional left projective space, thanks to the following:
@@
For a difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]$ which is not properly bounded, the kernel does not have to be free and/or of finite rank. Still, the kernel of such an operator is relatively small in the following sense:
@@
Let $\mathcal{D}\in R^{\mathbb{Z}}[T]$, $\mathcal{D}\neq 0$ be a difference operator of degree less than $d$. Then, for any properly bounded difference operator $\mathcal{D}^{\prime}\in R^{\mathbb{Z}}[T]$ of degree $d$, there exists $a\in\mathrm{Ker}\,\mathcal{D}^{\prime}$ such that $a\notin\mathrm{Ker}\,\mathcal{D}$.
@@
Since $\mathcal{D}\neq 0$, there is $i\in\mathbb{Z}$ such that the condition $\mathcal{D}a=0$, where $a\in R^{\mathbb{Z}}$, imposes a non-trivial relation on $d$ consecutive entries $a_{i},\dots,a_{i+d-1}$ of $a$. At the same time, the map $\mathrm{Ker}\,\mathcal{D}^{\prime}\to R^{d}$ given by $a\mapsto(a_{i},\dots,a_{i+d-1})$ is a bijection, so we can find an element of $\mathrm{Ker}\,\mathcal{D}^{\prime}$ whose entries do not satisfy the relation forced by $\mathcal{D}a=0$. ∎
Now, let $M$ be a free finite-rank $R$-module. Then $R^{\mathbb{Z}}\otimes_{R}M\simeq M^{\mathbb{Z}}$, as left $R^{\mathbb{Z}}$-modules (where $M^{\mathbb{Z}}$ consists of bi-infinite sequences valued in $M$). Any difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ acts on $M^{\mathbb{Z}}=R^{\mathbb{Z}}\otimes_{R}M$ by $\mathcal{D}(a\otimes x)=(\mathcal{D}a)\otimes x$, where $a\in R^{\mathbb{Z}}$, $x\in M$. Put differently, for $\mathcal{D}=\sum_{i=0}^{d}\alpha_{i}T^{i}$ and $v=(v_{i})_{i\in\mathbb{Z}}\in M^{\mathbb{Z}}$, we have
@@
$\left(\begin{array}[]{ccc}a_{1}&\dots&a_{m}\end{array}\right)=\left(\begin{array}[]{ccc}\vdots&&\vdots\\ v_{i,1}&\dots&v_{i,m}\\ \vdots&&\vdots\end{array}\right)$ (5.1)
@@
Suppose $\mathcal{D}\in R^{\mathbb{Z}}[T]$ is a difference operator, and $v\in M^{\mathbb{Z}}$, where $M$ is a free finite-rank left $R$-module. Then $\mathcal{D}v=0$ if and only if $\mathrm{Span}\,v\subset\mathrm{Ker}\,\,\mathcal{D}$.
@@
which is 0 if and only if $\mathcal{D}(a_{i}){\,=0}$ for all $i$, which is the same as to say $\mathrm{Span}\,v\subset\mathrm{Ker}\,\,\mathcal{D}$. ∎
The next proposition is one of the cornerstones of our construction and is where stable finiteness comes into play. For a sequence $v\in M^{\mathbb{Z}}$ such that $\mathcal{D}v=0$, it gives a criterion for when the associated sequence of points in $\mathbb{P}(M)$ is a polygon. This criterion is a difference counterpart of the following standard result from the theory of ordinary differential equations: for functions $f_{1}(x),\dots,f_{n}(x)$ belonging to a kernel of the differential operator
@@
Let $R$ be a stably finite ring. Suppose $\mathcal{D}\in R^{\mathbb{Z}}[T]$ is a properly bounded difference operator of degree $d$, and $v\in M^{\mathbb{Z}}$, where $M$ is a free left $R$-module of rank $d$. Suppose that $\mathcal{D}v=0$. Then the following conditions are equivalent:
@@
$V_{k}:=\left(\begin{array}[]{ccc}v_{k,1}&\dots&v_{k,d}\\ \vdots&&\vdots\\ v_{k+d-1,1}&\dots&v_{k+d-1,d}\end{array}\right)\in\mathrm{M}_{d}(R)$
where $v_{i,j}\in R$ are coordinates of $v_{i}$ relative to the basis $e_{1},\dots,e_{d}$. The rows of the matrix $V_{k}$ are coordinate expressions for $v_{k},\dots,v_{k+d-1}\in M$, so $V_{k}$ is left invertible if and only if $v_{k},\dots,v_{k+d-1}$ span $M$. At the same time, the matrix $V_{k}$ consists of $d$ consecutive rows of the infinite matrix (5.1) whose columns span the module $\mathrm{Span}\,v$. Since taking $d$ consecutive entries gives an isomorphism $\mathrm{Ker}\,\,\mathcal{D}\simeq R^{d}$, the columns of the matrix $V_{k}$ span $R^{d}$ if and only if the columns of the matrix (5.1) span $\mathrm{Ker}\,\,\mathcal{D}$, which is equivalent to saying that $\mathrm{Span}\,v=\mathrm{Ker}\,\,\mathcal{D}$. To sum up, $v_{k},\dots,v_{k+d-1}$ span $M$ if and only if the matrix $V_{k}$ is left-invertible, while $\mathrm{Span}\,v=\mathrm{Ker}\,\,\mathcal{D}$ if and only if $V_{k}$ is right-invertible. Since $R$ is stably finite, those conditions are equivalent. ∎
@@ 5.2 The correspondence between difference operators and polygons
Note that the ring $R^{\mathbb{Z}}[T]$ of difference operators over $R$ contains a subring $R^{\mathbb{Z}}$ of bi-infinite sequences valued in $R$: such sequences are just degree zero difference operators. In particular, the group of units $(R^{\mathbb{Z}})^{\times}$ in $R^{\mathbb{Z}}$ (which consists of bi-infinite sequences of units in $R$) acts on $R^{\mathbb{Z}}[T]$ in two different ways: by left multiplication, and by right multiplication.
@@
By the set of projective equivalence classes of polygons in the left projective space $\mathbb{P}^{d-1}(R)$, we mean the set of polygons in projectivizations of all rank $d$ free left $R$-modules, considered up to projective equivalence.
Similarly, there is a natural one-to-one correspondence between the set of projective equivalence classes of polygons in the right projective space $\mathbb{P}^{d-1}(R)$, and the set of properly bounded right difference operators of degree $d$ with coefficients in $R$, up to multiplication on the left and right by bi-infinite sequences of units in $R$.
In the case when the base ring $R$ is a field, the correspondence between polygons and difference operators is well-known. In particular, the case $R=\mathbb{R}$ follows from [9, Proposition 3.3].
In future sections, we are most interested in the case of $d=3$, i.e., a correspondence between polygons in the projective plane $\mathbb{P}^{2}(R)$ and degree $3$ difference operators. Still, we give a proof in the general case because it works in the same way for any $d$. Also, we expect $d>3$ cases of Theorem 5.7 to be useful for the study of pentagram maps in higher-dimensional projective spaces, cf. [9].
@@
Let us denote the space of all degree $d$ properly bounded difference operators over $R$ as $R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}$ and the space of polygons in $\mathbb{P}^{d-1}(R)$ up to projective equivalence as $\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Our first aim to build a map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
Let $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}$. Then, by Proposition 5.3, we know that $\mathrm{Ker}\,\,\mathcal{D}\subset R^{\mathbb{Z}}$ is a rank $d$ free right $R$-module. Thus, $(\mathrm{Ker}\,{\mathcal{D}})^{*}=\mathrm{Hom}_{R}(\mathrm{Ker}\,{\mathcal{D}},R)$ is a free rank $d$ left $R$-module. For any $i\in\mathbb{Z}$, let $\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,{\mathcal{D}})^{*}$ be defined by $\varepsilon_{\mathcal{D},i}(a):=a_{i}$ for any $a\in\mathrm{Ker}\,{\mathcal{D}}$. Then $\varepsilon_{\mathcal{D}}\in((\mathrm{Ker}\,{\mathcal{D}})^{*})^{\mathbb{Z}}=R^{\mathbb{Z}}\otimes_{R}(\mathrm{Ker}\,{\mathcal{D}})^{*}$.
@@
For stably finite ring $R$, we have $\mathrm{Span}\,\varepsilon_{\mathcal{D}}=\mathrm{Ker}\,\mathcal{D}$. In particular, $\mathcal{D}\varepsilon_{\mathcal{D}}=0$ (recall that a difference operator $\mathcal{D}=\sum_{i=0}^{d}\alpha_{i}T^{i}$ acts on $\varepsilon_{\mathcal{D}}$ by $(\mathcal{D}\varepsilon_{\mathcal{D}})_{j}=\sum_{i=0}^{d}\alpha_{ij}\varepsilon_{\mathcal{D},i+j}$).
@@
For stably finite ring $R$ and for any $i\in\mathbb{Z}$, the list $\varepsilon_{\mathcal{D},i},\dots,\varepsilon_{\mathcal{D},i+d-1}$ is a basis of $(\mathrm{Ker}\,\,\mathcal{D})^{*}$.
@@
It follows that the sequence of points $P_{i}=R\varepsilon_{\mathcal{D},i}$ is a polygon in $\mathbb{P}((\mathrm{Ker}\,\,\mathcal{D})^{*})$. Thus, every properly bounded difference operator gives rise to a polygon in the projectivization of the dual to its kernel. This gives a map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Denote the collection of properly bounded, degree $d$ left difference operators up to multiplication on the left and right by bi-infinite sequences of units as ${R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})$.
@@
For stably finite ring $R$, the map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ factors through a map $\varphi:{R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
@@
We need to check that if we take a properly bounded left difference operator $\mathcal{D}$, and multiply it on the left or right by a bi-infinite sequence of units in $R$, then the image of the resulting operator under $\bar{\varphi}$ is in the same projective equivalence class as the image of $\mathcal{D}$. Consider $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}$ and a bi-infinite sequence $c\in(R^{\mathbb{Z}})^{\times}$ of units in $R$. Note that multiplication on the left $\mathcal{D}\mapsto c\mathcal{D}$ does not change the kernel and hence the image under $\bar{\varphi}$. So, it suffices to establish projective equivalence of the polygons $\bar{\varphi}(\mathcal{D}c)$ and $\bar{\varphi}(\mathcal{D})$. Clearly, left multiplication by $c$ gives a right $R$-module isomorphism $\lambda_{c}\colon\mathrm{Ker}\,\mathcal{D}c\to\mathrm{Ker}\,\mathcal{D}$. Consider the associated left module isomorphism $\lambda_{c}^{*}\colon(\mathrm{Ker}\,\mathcal{D}c)^{*}\to(\mathrm{Ker}\,\mathcal{D})^{*}$. As before, let $\varepsilon_{\mathcal{D}c,i}\in(\mathrm{Ker}\,\mathcal{D}c)^{*}$ be defined by $\varepsilon_{\mathcal{D}c,i}(a)=a_{i}$, and let $a\in\mathrm{Ker}\,\mathcal{D}$. Then
@@
Let $R$ be a stably finite ring. Suppose that $\mathcal{D},\mathcal{D}^{\prime}\in R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}$ are properly bounded (left) difference operators of degree $d$. Assume also that $\mathrm{Ker}\,\mathcal{D}=\mathrm{Ker}\,\mathcal{D}^{\prime}$. Then $\mathcal{D}^{\prime}=c^{\prime}\mathcal{D}$ for some $c^{\prime}\in(R^{\mathbb{Z}})^{\times}$.
@@
Since $\mathcal{D}^{\prime}$ and $\mathcal{D}$ are properly bounded of the same degree $d$, there is a sequence of units $c^{\prime}\in(R^{\mathbb{Z}})^{\times}$ such that $\mathcal{R}:=\mathcal{D}^{\prime}-c^{\prime}\mathcal{D}$ has degree at most d1d-1. Note that since $\mathrm{Ker}\,\mathcal{D}^{\prime}=\mathrm{Ker}\,\mathcal{D}$, we have $\mathrm{Ker}\,\mathcal{R}\supset\mathrm{Ker}\,\mathcal{D}$. So, by Proposition 5.4, it follows that $\mathcal{R}=0$, meaning that $\mathcal{D}^{\prime}=c^{\prime}\mathcal{D}$, as desired. ∎
@@
For stably finite ring $R$, the map $\varphi:{R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ is injective.
@@
Consider two properly bounded left difference operators $\mathcal{D},\mathcal{D}^{\prime}\in R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}$ such that $\overline{\varphi}(\mathcal{D})=\overline{\varphi}(\mathcal{D}^{\prime})$. We need to show that $\mathcal{D}$ and $\mathcal{D}^{\prime}$ belong to the same class in the quotient ${R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}}/((R^{\mathbb{Z}})^{\times}\times(R^{\mathbb{Z}})^{\times})$.
@@
Now, to complete the proof of Theorem 5.7, it suffices to show that the mapping $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ is surjective. In other words, given a polygon $(P_{i})$ in $\mathbb{P}(M)$, where $M$ is a rank $d$ free left $R$-module, we need to find a difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{\mathrm{pb}}$ such that the associated polygon in $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})$ is projectively equivalent to $(P_{i})$.